### v1

@@ 6 Conclusions
This study proposed a two-step procedure for assessing the level of appropriateness of Italian referrals. We presented a case study on the venous echocolordoppler of the lower limbs in the Lombardy Region, demonstrating excellent performance on a validation set and extracting relevant insights for public health authorities. The outcomes of this analysis enabled the Lombardy Region to reinforce existing guidelines for physicians, aiming to reduce the number of inappropriate referrals. By addressing one of the key drivers of inefficiency in the healthcare system, this approach has the potential to alleviate significant burdens on the Italian NHS, lowering waiting times and enhancing the overall effectiveness of healthcare services.
@@
  3. 3.
    GP: General Practitioner
  4. 4.
    EHR: Electronic Health Record
  5. 5.
    ATS/LHA: Agenzia di Tutela della Salute / Local Health Authority
  6. 6.
    GDPR: General Data Protection Regulation
  7. 7.
    IQR: Interquartile Range
  8. 8.
    ICD: International Classification of Diseases
  9. 9.
    BERT: Bidirectional Encoder Representation for Transformers
  10. 10.
    DGR: Decreto della Giunta Regionale (Regional Law)
  11. 11.
@@
  12. 12.
    CQ: Clinical Question
@@
VT: investigation, formal analysis, methodology, software, data curation, writing - original draft AB: resources, data curation, conceptualization ME: resources, data curation, conceptualization OL: resources, data curation, conceptualization FI: conceptualization, methodology, formal analysis, supervision, writing - review & editing
Acknowledgments
@@ Appendix A Manual annotations
A validation set of 1000 clinical questions was manually annotated for the main disease/symptom determining the referrals. The existing guidelines from the Lombardy Region [26] were used as a reference, but labels were not forced to be one of the categories mentioned in the guidelines as appropriate or inappropriate. If none of the categories mentioned in the guidelines fit the reason for the referral, a new category representing the main disease or symptom mentioned in the referral was created. The other category was used only when it was impossible to identify a symptom or a disease determining the referral (e.g., “in-depth examination to actually establish whether there is an ongoing pathology given the symptoms present” or “gfwhdfg”).
@@
This procedure, after four iterations, including consultation with an angiologist, defined the labels reported in Table A.1.
Table A.1: Labels derived by manual annotations on the validation set of 1000 clinical questions for venous echocolordoppler of the lower limbs. CQ = Clinical Questions, REF = Referrals.
@@
Appendix B Clustering
B.1 Additional details on clustering procedures
For TF-IDF representation, we limited the vocabulary to the top 1,000 most frequent tokens. Clustering was performed with HDBscan and K-Means, obtaining the best results with K-Means. The number of clusters was selected with the knee-elbow method analyzing WSS with k from 2 to 50, which led to k=40.
For LDA, the number of topics (clusters) was selected by analyzing coherence with k from 2 to 50, which led to k=7.
For HDBscan, the initial minimum size was set to 500, selected as the minimum size that led to the emergence of at least 5 clusters, and was progressively halved up to 20.
B.2 Additional clustering results
Table LABEL:tab:clusters-summ presents all the clusters identified for venous echocolordoppler of the lower limbs, with their keyword descriptions, the sizes and the labels they were assigned.
Table B.1: Summary of all clusters identified for echocolordoppler of the lower limbs, including their keyword representation, the label assigned them and the number of distinct clinical questions (CQ) and referrals (REF) that they contain.
CLUSTER KEYWORDS [ITA] CLUSTER KEYWORDS [ENG] CLUSTER LABEL # CQ # REF
claudicatio; riferita claudication; referred claudication 1132 1686
claudicatio; intermittens claudicatio; intermittens claudication 653 1179
claudicatio claudicatio claudicatio 402 530
claudicatio; sospetta; claudicazio claudicatio; suspected; claudication claudicatio 60 126
claudicatio; arti; inferiori claudicatio; limbs; lower claudicatio 41 49
dermatite dermatitis dermatitis 421 474
diabete diabetes diabetes 1540 3249
diabete; mellito diabetes; mellitus diabetes 534 1847
diabete; cardiopatia; mellito; diabetes; heart disease; mellitus diabetes 755 1254
diabetico; diabetica diabetes; diabetic diabetes 209 165
dislipidemia; diabete dyslipidaemia; diabetes diabetes 36 68
dialisi dialysis dialysis 110 222
dimero dimer dimer 1343 1601
dispnea dyspnoea dyspnoea 104 171
dolore pain pain 6343 12051
deambulazione; dolore; difficolta walking; pain; difficulty pain 1294 3884
algia; artralgia; piedi algia; arthralgia; foot pain 1028 2491
polpaccio; dolore calf; pain pain 707 1753
algie; arti; gambe; polpacci; bilaterali; natura; determinare pain; limbs; calves; bilateral; nature; determine pain 160 631
dolore; natura; determinare; algie pain; nature; determine; pain pain 53 126
dolore; popliteo pain; popliteal pain 35 117
algia; covid; gamba pain; covid; leg pain 30 63
dolore; polpaccio; natura; determinare pain; calf; nature; determine pain 12 10
crampi; notturni cramps; nocturnal pain 6 5
edema edema edema 11686 20723
edemi; declivi edema; declivities edema 9491 16473
edema; arti; edemi edema; limbs; edemas edema 3712 7867
arti; inferiori; edemi limbs; lower; edemas edema 528 2486
edema; gamba; piede; polpaccio edema; leg; foot; calf edema 1046 2467
edema; arti; inferiori edema; limbs; lower edema 899 2350
edema; perimalleolare edema; perimalleolar edema 1039 2214
edema; natura; determinare edema; nature; determine edema 1027 2199
edemi; perimalleolari edema; perimalleolar edema 1190 2151
edema; malleolare edema; malleolar edema 935 1136
edemi; vasculopatia; declivi; edema edema; vasculopathy; declivities edema 529 602
edema; flebite edema; phlebitis edema 217 297
arti; bilaterale; piede; edema; succulenza; perimalleolare limbs; bilateral; foot; edema; succulence; perimalleolar edema 117 192
arti; inferiori; edema limbs; lower; edema edema 80 168
arti; inferiori; natura; determinare; edema; edemi limbs; lower; nature; determine; edema edema 80 168
edema; inferiori; arti edema; lower; limbs edema 74 163
edema; sinistra edema; left edema 61 108
claudicatio; edemi; edema; cardiopatia; arti; inferiore; lieve claudication; edema; edema; cardiopathy; limbs; lower; mild edema 32 34
linfedema; arti lymphedema; limbs edema 23 24
emorragia haemorrhage haemorrhage 18 29
formicolio tingling tingling 43 170
freddo cold cold 49 147
fumatore smoker smoker 121 193
gravidanza pregnancy pregnancy 92 225
insufficienza respiratoria respiratory insufficiency respiratory insufficiency 724 1060
venosa; insufficienza venous; insufficiency venous insufficiency 30324 71150
venosa; arti; inferiori; insufficienza; sospetta venous; limbs; lower; insufficiency; suspected venous insufficiency 19532 28831
superficiale; venosa; insufficienza superficial; venous; insufficiency venous insufficiency 276 613
insufficienza; venosa; cronica venous; insufficiency; chronic venous insufficiency 415 506
venosa; safenectomia; insufficienza venous; saphenectomy; insufficiency venous insufficiency 280 311
venosa; claudicatio; insufficienza venous; claudication; insufficiency venous insufficiency 109 180
venosa; arti; insufficienza; varici venous; limb; insufficiency; varicose venous insufficiency 110 112
insufficienza; venosa; arti; inferiori venous; insufficiency; limbs; lower venous insufficiency 23 102
venoso; circolo; deficit; arti; venous; circulation; insufficiency; limbs venous insufficiency 74 102
venosa; insufficienza; arti; inferiori; edemi venous; insufficiency; limbs; lower; edemas venous insufficiency 47 58
insufficienza; venosa venous; insufficiency venous insufficiency 0 0
arti; venosa; inferiori limbs; venous; lower venous insufficiency 0 0
venosa; edemi; declivi; insufficienza venous; edemas; declivity; insufficiency venous insufficiency 0 0
venosa; insufficienza; edemi; arti; inferiori venous; insufficiency; edemas; limbs; lower venous insufficiency 0 0
venosa; cronica venous; chronic venous insufficiency 0 0
insufficienza; venosa; cronica; sospetta venous; insufficiency; chronic; suspected venous insufficiency 264 374
intervento; ginocchio; artroprotesi; valutazione surgery; knee; arthroplasty; evaluation intervention 67 77
lista; trapianto; rene; inserimento list; transplant; kidney; insertion intervention 61 69
lista; trapianto; valutazione; inserimento; rene; sinistra; dialisi; intervento list; transplant; evaluation; insertion; kidney; left; dialysis; intervention intervention 49 52
ipercolesterolemia hypercholesterolaemia hypercholesterolaemia 363 416
ipoestesia hypoesthesia hypoesthesia 31 80
ischemia ischaemia ischaemia 517 617
parestesie paresthesias paresthesias 1834 3965
parestesie; gamba; sinistra paresthesias; leg; left paresthesias 541 891
pesantezza heaviness heaviness 215 475
arti; inferiori; linfedema; pesantezza limbs; lower; lymphedema; heaviness heaviness 25 124
arti; inferiori; pesantezza limbs; lower; heaviness heaviness 18 106
safenectomia saphenectomy saphenectomy 1988 4177
safenectomia; esiti saphenectomy; outcomes saphenectomy 1374 3617
safenectomia; pregressa saphenectomy; previous saphenectomy 342 379
safenectomia; edemi saphenectomy; edemas saphenectomy 41 117
sindrome metabolica metabolic syndrome metabolic syndrome 12 48
stent stent stent 301 612
tromboflebite thrombophlebitis thrombophlebitis 7073 17415
flebite; sospetta phlebitis; suspected thrombophlebitis 5408 12379
flebite; gamba; superficiale phlebitis; leg; superficial thrombophlebitis 3037 8850
varicoflebite; sospetta; tromboflebite varicophlebitis; suspected; thrombophlebitis thrombophlebitis 4546 7739
tromboflebite; superficiale thrombophlebitis; superficial thrombophlebitis 950 1648
gamba; tromboflebite; destra leg; thrombophlebitis; right thrombophlebitis 397 486
tromboflebite; safena thrombophlebitis; saphenous vein thrombophlebitis 386 384
flebite; coscia; superficiale phlebitis; thigh; superficial thrombophlebitis 109 194
flebopatia; sospetta; arti; inferiori; sospetta phlebopathy; suspected; lower; limbs thrombophlebitis 209 175
tromboflebite; poplitea; arti; inferiori thrombophlebitis; popliteal; lower; limbs thrombophlebitis 54 102
flebite; superficiale; gamba phlebitis; superficial; leg thrombophlebitis 48 97
tromboflebite; sospetta; thrombophlebitis; suspected thrombophlebitis 30 68
trombosi thrombosis thrombosis 17060 43558
trombosi; venosa; profonda thrombosis; venous; deep thrombosis 17453 38230
trombosi; vascolare; thrombosis; vascular thrombosis 5086 10180
trombosi; venosa venous; thrombosis thrombosis 1325 3337
venosa; trombosi; profonda; sospetto; segni; clinici; clinico venous; thrombosis; deep; suspected; signs; clinical thrombosis 566 1152
poplitea; trombosi popliteal; thrombosis thrombosis 607 642
venosa; trombosi; poplitea venous; thrombosis; popliteal thrombosis 407 525
trombosi; safena; insufficienza thrombosis; saphenous; insufficiency thrombosis 362 472
segni; clinici; sospetto; clinico; trombosi; venosa; profonda; edema signs; clinical; suspected; clinical; venous; thrombosis; deep; edema thrombosis 342 389
safena; trombosi saphenous vein; thrombosis thrombosis 301 263
post; trombotica post; thrombotic thrombosis 36 73
tumefazione swelling swelling 533 741
tumefazione; gamba swelling; leg swelling 172 448
destro; polpaccio; tumefazione; paziente right; calf; swelling; patient swelling 35 49
ulcera; vascolare ulcer; vascular ulcer 828 3861
ulcera ulcer ulcer 1258 2411
varici varices varices 13999 35939
arti; inferiori; varici limbs; lower; varices varices 7018 18157
venose; varici venous; varices varices 4957 13554
varicosita; stasi varicosities; stasis varices 957 2701
varici; arti; inferiori varices; limbs; lower limbs varices 219 277
vene; varicose; algie veins; varicose; algie varices 121 194
safena; varici saphenous veins; varicose varices 47 136
varicosita; varicosita’ varicosities; varicositis varices 86 97
vascolari; esiti; polmonite; varici; chirurgici; linfangite vascular; outcomes; pneumonia; varices; surgical; lymphangitis varices 60 88
varici; sospette varicose veins; suspected varices 34 63
varicosita; gamba varicosities; leg varices 80 63
varici; intervento varicosities; surgery varices intervention 3144 7126
vasculopatia; periferica vasculopathy; peripheral vasculopathy 3820 8200
vasculopatia vasculopathy; peripheral vasculopathy 769 1886
vasculopatia; diabetico vasculopathy; diabetic vasculopathy 393 408
arteriopatia; obliterante arteriopathy; obliterative vasculopathy 288 272
arteriopatia; obliterante; periferica arteriopathy; obliterative; peripheral vasculopathy 47 117
iperteso; vasculopatico; vasculopatia; dislipidemico hypertensive; vasculopathy; dyslipidemic vasculopathy 17 97
aneurisma; aorta; esiti; popliteo; addominale; aortico aneurysm; aorta; popliteal; abdominal; aortic vasculopathy 42 78
vasculopatia; arti; inferiori vasculopathy; limbs; lower vasculopathy 36 78
vasculopatia; arti; inferiori; polidistrettuale vasculopathy; limb; lower; polydistrict vasculopathy 35 49
vasculopatia; venosa vasculopathy; venous vasculopathy 25 34
arteriopatia arteriopathy vasculopathy 20 25
natura; determinare; periferica nature; determine; peripheral other 3978 10039
natura; determinare; arti; inferiori nature; determine; limbs; lower other 5018 9601
paziente; polsi; periferici patient; wrists; peripheral other 2018 8015
arti; inferiori limbs; lower other 2190 5165
safena; incontinenza saphenous; incontinence other 288 782
ipertrofia hypertrophy other 506 591
tromboembolia; polmonare; esiti thromboembolism; pulmonary; outcomes other 386 379
embolia; polmonare embolism; pulmonary other 110 174
femoro; femorale; paziente; popliteo; sospetta; asse; esiti; venosa femoral; femoral; patient; popliteal; suspected; axis; outcomes; venous other 49 49
stenosi; venoso stenosis; venous other 30 49
dislipidemia; ipertensione dyslipidaemia; hypertension other 25 44
aorta; sottorenale aorta; subrenal other 25 29
Table B.1: *
Appendix C Pre-processing results
Table C.1 reports the list of the 50 most frequent abbreviations that were expanded by our pre-processing, while Table C.2 contains an example list of 50 misspelt words that were detected and corrected during our pre-processing.
Table C.1: Most frequent abbreviations in clinical questions for echocolordoppler of the venous limbs, along with their expansions in Italian and English.
@@ Appendix C Pre-processing results
Table C.2: Sample of 50 misspelt words with the corresponding replacements applied by our pre-processing.

### v2

@@ 5 Conclusions
This study proposed a methodological pipeline for assessing the level of appropriateness of Italian referrals. We presented two case studies on venous echocolordoppler of the lower limbs and flexible endoscope colonoscopy in the Lombardy Region, demonstrating excellent performance on manually annotated test sets and deriving relevant insights for public health authorities. The outcome of this analysis allowed the Lombardy Region to reinforce existing guidelines for physicians, with the aim of reducing the number of inappropriate referrals. This work is an example of how artificial intelligence can be deployed on large, real-world administrative datasets, highlighting the potential of AI to support healthcare systems in monitoring practices, informing policy, and ultimately alleviating significant burdens on the Italian NHS by lowering waiting times and enhancing overall effectiveness of care.
Format: pdf Title: Supplementary Material Description: contains more details about the manual annotation process, hyperparameters selection and additional results
@@
    AI: Artificial Intelligence
  2. 2.
    BERT: Bidirectional Encoder Representation for Transformers
  3. 3.
    CIBD: Chronic Inflammatory Bowel Disease
  4. 4.
    CQ: Clinical Question
  5. 5.
    DGR: Decreto della Giunta Regionale (Regional Law)
  6. 6.
    ECD: venous EcoColorDoppler of the lower limbs
  7. 7.
    EHR: Electronic Health Record
  8. 8.
    FEC: Flexible Endoscope Colonscopy with possible biopsies
  9. 9.
    GP: General Practitioner
  10. 10.
    ICD: International Classification of Diseases
  11. 11.
    IQR: InterQuartile Range
  12. 12.
    LDA: Latent Dirichlet Allocation
  13. 13.
    ATS/LHA: Agenzia di Tutela della Salute / Local Health Authority
  14. 14.
@@
  16. 16.
@@
  17. 17.
    TF-IDF: Term Frequency - Inverse Document Frequency
@@
VT: analysis and interpretation of data, development of the methodology, drafting of the manuscript AB: concept and design of the study, acquisition, analysis and interpretation of data ME: concept and design of the study, acquisition, analysis and interpretation of data OL: concept and design of the study, acquisition, analysis and interpretation of data FI: concept and design of the study, analysis and interpretation of data, critical revision of the manuscript, supervision. All authors read and approved the final manuscript.
Acknowledgements
@@ Appendix A Manual annotations
A validation set of 1,000 clinical questions was manually annotated for the main disease/symptom determining the referrals, for both ECD and FEC datasets. The existing guidelines from the Lombardy Region [31, 32] were used as a reference, but labels were not forced to be one of the categories mentioned in the guidelines as appropriate or inappropriate. If none of the categories mentioned in the guidelines fit the reason for the referral, a new category representing the main disease or symptom mentioned in the referral was created. The other category was used only when it was impossible to identify a symptom or a disease determining the referral (e.g., “in-depth examination to actually establish whether there is an ongoing pathology given the symptoms present” or “gfwhdfg”).
@@
This procedure, after four iterations, including consultation with specialist physicians, defined the labels reported in Tables A.1 and A.2. After review of these labels, some of them were merged:
  • bleeding and occult blood since bleeding in this context is always an occult bleeding
  • colitis, chron’s disease, ulcerative rectocolitis, chronic inflammatory bowel disease, under the latter which is a category comprehensive of all these diseases
Table A.1: Labels derived by manual annotations on the validation set of 1,000 clinical questions for ECD.
@@
Table A.2: Labels derived by manual annotations on the validation set of 1,000 clinical questions for FEC
Label # CQ # REF
Polyps 158 344
Bleeding 127 959
Colon cancer 107 187
Abdominalgia 104 214
Colon cancer familiarity 82 141
Anaemia 65 115
Prevention 57 161
Occult blood 36 789
Constipation 36 139
Diverticulosis 33 48
Altered bowel 30 65
Colitis 28 40
Chronic diarrhoea 28 107
Diverticulitis 14 28
Crohn’s disease 12 16
Haemorrhoidal disease 9 30
Chronic inflammatory bowel disease 8 15
Ulcerative rectocolitis 8 15
Surgery 5 6
Intestinal obstruction 5 10
Dyspepsia 4 5
Weight loss 3 8
Irritable bowel syndrome 3 6
Lymphoma 3 4
Transplant 3 15
Other 2 2
Colonpathy 2 7
Enteritis 2 3
Proctitis 2 4
Rectal prolapse 2 12
Lynch syndrome 2 2
Angiodysplasia 1 1
Appendicitis 1 1
Celiac disease 1 1
Intestinal loops distension 1 1
Hepatopathy 1 1
Protruding formation 1 1
Gastritis 1 2
Ileitis 1 1
Hypercholesterolaemia 1 1
Splenic flexure thickening 1 3
Rectal wall thickening 1 2
Sigma wall thickening 1 1
Behcet’s disease 1 1
Liver disease 1 1
Pancreatitis 1 1
Proctalgia 1 3
Sigmoiditis 1 1
Suspected colic lesion 1 33
Splenic fissure stenosis 1 1
Rectal tenesmus 1 2
Appendix B Pre-processing results
Table B.1 reports the list of the 50 most frequent abbreviations that were expanded by our pre-processing, while Table B.2 contains an example list of 50 misspelt words that were detected and corrected during our pre-processing. Table B.3 reports the summary results on disease labels without applying pre-processing.
Table B.1: Most frequent abbreviations in clinical questions for echocolordoppler of the venous limbs, along with their expansions in Italian and English.
@@ Appendix B Pre-processing results
Table B.2: Sample of 50 misspelt words with the corresponding replacements applied by our pre-processing.